### v2

# Fully Adaptive Zeroth-Order Method for Minimizing Functions with Compressible Gradients

Abstract
We propose an adaptive zeroth-order method for minimizing differentiable functions with $L$-Lipschitz continuous gradients. The method is designed to take advantage of the eventual compressibility of the gradient of the objective function, but it does not require knowledge of the approximate sparsity level $s$ or the Lipschitz constant $L$ of the gradient. We show that the new method performs no more than $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of an objective function with $n$ variables. Assuming additionally that the gradients of the objective function are compressible, we obtain an improved complexity bound of $\mathcal{O}\left(s\log\left(n\right)\epsilon^{-2}\right)$ function evaluations, which holds with high probability. Preliminary numerical results illustrate the efficiency of the proposed method and demonstrate that it can significantly outperform its non-adaptive counterpart. Accompanying code can be found at https://github.com/mines-opt-ml/ZORO-FA.
Keywords: Derivative-Free Optimization; Black-Box Optimization; Zeroth-Order Optimization; Worst-Case Complexity; Compressible Gradients
1 Introduction
In this work we consider the unconstrained minimization of a possibly nonconvex differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}$. Most well-known iterative algorithms for this type of problem are gradient-based, meaning at each iteration they require at least one evaluation of the gradient of $f(\,\cdot\,)$. However, in many problems, it is impossible to compute the gradient of the underlying objective function. Examples include simulation-based optimization [18], reinforcement learning [22, 15, 8] and hyperparameter tuning [3, 25]. These settings create a demand for zeroth-order optimization algorithms, which are algorithms that rely only on evaluations of the objective function $f(\,\cdot\,)$. Recently, problems requiring zeroth-order methods have emerged in which $n$, the dimension of the problem, is of order $10^{5}$ or greater. In particular, we highlight many recent works applying zeroth-order methods to fine-tune Large Language Models (LLMs) [14, 28, 13, 7]. For problems of this scale, classical zeroth-order algorithms can struggle to make progress as they generically require $\mathcal{O}(n)$ function evaluations (“queries”) per iteration. As such, various structural assumptions on $\nabla f(x)$ have recently been proposed which allow for robust algorithms making as few as $\mathcal{O}(\log n)$ queries per iteration with high probability.
Here, we focus on methods that exploit (approximate) gradient sparsity [26, 1, 4, 5, 6, 29], particularly the ZORO-type algorithms [4, 6] which use tools from signal processing to robustly approximate $\nabla f(x)$. One adverse consequence of this fusion of optimization and signal processing techniques is that ZORO-type algorithms have an unusually large number of hyperparameters. In this paper, we address this issue by proposing a fully adaptive version of ZORO, which we call ZORO-FA. The proposed method is designed to take advantage of the eventual compressibility of the gradient of the objective function, in which case a suitable gradient approximation can be obtained with high probability using $\mathcal{O}\left(s\log(n)\right)$ function evaluations. Specifically, at each iteration, the new method first attempts to use these cheap gradient approximations, and only when they fail is a forward-finite difference gradient computed using $\mathcal{O}\left(n\right)$ function evaluations. Moreover, ZORO-FA is endowed with a novel line-search procedure for selecting multiple hyperparameters at once, adapted from that proposed in [11]. We show that the new method takes at most $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find $\epsilon$-approximate stationary points. Assuming that the gradients of the objective function are compressible with sparsity level $s$, we obtain an improved complexity bound of $\mathcal{O}\left(s\log\left(n\right)\epsilon^{-2}\right)$ function evaluations, which holds with high probability. Our preliminary numerical results indicate that ZORO-FA requires significantly fewer function evaluations than ZORO to find points with similar function values, and that the improvement factor increases as the dimensionality of the problems grows.
Contents. The rest of the paper is laid out as follows. In Section 2 we present the key auxiliary results about our gradient approximation and the corresponding variant of the descent lemma. In Section 3, we present the new zeroth-order method and establish a worst-case oracle complexity bound. Finally, in Section 4 we report some numerical results.
2 Auxiliary Results
In what follows we will consider the following assumptions:
  • A1
    $\nabla f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $L$-Lipschitz continuous.
  • A2
    $f$ is bounded below, that is there exists $f_{\mathrm{low}}\in\mathbb{R}$ such that
    $f(x)\geq f_{\mathrm{low}}\quad\text{ for all }x\in\mathbb{R}^{n}.$
  • A3
    There exists $p\in(0,1)$ such that, for all $x\in\mathbb{R}^{n}$, $\nabla f(x)$ is $p$-compressible with magnitude $\|\nabla f(x)\|$, i.e.,
    $\left|\nabla f(x)\right|_{(j)}\leq\dfrac{\|\nabla f(x)\|}{j^{1/p}},\quad j=1,\ldots,n,$
    where $|u|_{(j)}$ denotes the $i$-th largest-in-magnitude component of $u\in\mathbb{R}^{n}$.
Given $x\in\mathbb{R}^{n}$ and $s\in\left\{1,\ldots,n\right\}$, let
$\left[\nabla f(x)\right]_{(s)}=\arg\min\left\{\|v-\nabla f(x)\|_{2}\,:\,\|v\|_{0}\leq s\right\}.$
Notice that $\left[\nabla f(x)\right]_{(s)}$ is the $s$-sparse vector formed by the $s$ largest (in modulus) entries of $\nabla f(x)$.
Lemma 2.1.
Suppose that A3 holds and let $s\in\left\{1,\ldots,n\right\}$. Then, for all $x\in\mathbb{R}^{n}$ we have
$\|\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\|_{1}\leq\dfrac{\|\nabla f(x)\|_{2}}{\left(\frac{1}{p}-1\right)}s^{1-1/p},$
and
$\|\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\|_{2}\leq\dfrac{\|\nabla f(x)\|_{2}}{\left(\frac{2}{p}-1\right)^{\frac{1}{2}}}s^{1/2-1/p}.$
Proof.
See Section 2.5 in [19].. ∎
We say that a matrix $Z\in\mathbb{R}^{m\times n}$ has the $4s$-Restricted Isometry Property ($4s$-RIP) when, for every $v\in\mathbb{R}^{n}$ with $\|v\|_{0}\leq 4s$ we have
$\left(1-\delta_{4s}(Z)\right)\|v\|_{2}^{2}\leq\|Zv\|_{2}^{2}\leq\left(1+\delta_{4s}(Z)\right)\|v\|_{2}^{2}$
for some constant $\delta_{4s}(Z)\in(0,1)$.
Lemma 2.2.
(Proposition 3.5 in [19]) If $Z\in\mathbb{R}^{m\times n}$ satisfies the $4s$-RIP for some $s\in\left\{1,\ldots,n\right\}$, then for any $v\in\mathbb{R}^{n}$ we have
$\|Zv\|_{2}\leq\sqrt{1+\delta_{4s}(Z)}\left(\|v\|_{2}+\frac{1}{\sqrt{s}}\|v\|_{1}\right).$
Consider the function $c_{0}:\mathbb{R}\to\mathbb{R}$ given by
$c_{0}(a)=\dfrac{a^{2}}{4}-\frac{a^{3}}{6}.$ (1)
Lemma 2.3.
(Theorem 5.2 in [2]) Given $s\in\mathbb{N}$ such that $m:=\lceil bs\ln(n)\rceil<n$, let $Z\in\mathbb{R}^{m\times n}$ be defined as
$Z=\frac{1}{\sqrt{m}}\left[z_{1}\ldots z_{m}\right]^{T},$
where $z_{i}\in\mathbb{R}^{n}$ are Rademacher random vectors [^1] That is, $[z_{i}]_{j}=\pm 1$ with equal probability.. Given $\delta\in(0,1)$, if
$b>\dfrac{4\left(1+\dfrac{1+\ln(\frac{12}{\delta})}{\ln(\frac{n}{4s})}\right)}{c_{0}(\frac{\delta}{2})},$ (2)
for $c_{0}(\,\cdot\,)$ defined in (1), then $Z$ satisfies the $4s$-RIP for $\delta_{4s}(Z)=\delta$, with probability $\geq 1-2e^{-\gamma m}$, where
$\gamma=\left[c_{0}\left(\frac{\delta}{2}\right)-\frac{4}{b}\left(1+\dfrac{1+\ln(\frac{12}{\delta})}{\ln(\frac{n}{4s})}\right)\right].$
2.1 Gradient Approximation
ZORO [6] uses finite differences along random directions to approximate directional derivatives, thought of as linear measurements of $\nabla f(x)$. More precisely, let $\left\{z_{i}\right\}_{i=1}^{m}\subset\mathbb{R}^{n}$ be Rademacher random vectors. For fixed $h>0$ define the vector $y\in\mathbb{R}^{m}$ with
$y_{i}=\dfrac{f(x+hz_{i})-f(x)}{h\sqrt{m}},\quad j=i,\ldots,m.$ (3)
and the matrix
$Z=\dfrac{1}{\sqrt{m}}\left[z_{1}\ldots z_{m}\right]^{T}\in\mathbb{R}^{m\times n}.$ (4)
Lemma 2.4.
Suppose that A1 holds and let $y$ and $Z$ be defined by (3) and (4), respectively. Then
$y=Z\nabla f(x)+hw,$ (5)
where $\|w\|\leq\frac{Ln}{2}$.
Proof.
From A1,
$\left|f(x+hz_{i})-f(x)-\nabla f(x)^{T}(hz_{i})\right|$ $\leq\frac{h^{2}}{2}L\|z_{i}\|_{2}^{2}$
$=\frac{Lh^{2}}{2}n\quad\text{ as $[z_{i}]_{j}=\pm 1$ for all $j$}$
Then
$\left|\frac{1}{\sqrt{m}}\dfrac{f(x+hz_{i})-f(x)}{h}-\frac{1}{\sqrt{m}}\nabla f(x)^{T}z_{i}\right|\leq\frac{Ln}{2\sqrt{m}}h.$
Now, writing $y$ as in (5), it follows from (4) that
$h|\left[w\right]_{i}|=|\left[y\right]_{i}-\left[Z\nabla f(x)\right]_{i}|\leq\frac{Ln}{2\sqrt{m}}h.$
Therefore,
$|\left[w\right]_{i}|\leq\dfrac{Ln}{2\sqrt{m}},\quad i=1,\ldots,m,$
which implies that $\|w\|_{2}\leq\frac{Ln}{2}$. ∎
We view (5) as a linear inverse problem to be solved for $\nabla f(x)$. As in [6], given $s\in\mathbb{N}\setminus\left\{0\right\}$, we will compute an approximation to $\nabla f(x)$ by approximately solving the problem
$\min\|Zv-y\|\quad\text{s.t.}\quad\|v\|_{0}\leq s,$ (6)
using the algorithm CoSaMP proposed in [19].
Lemma 2.5.
Suppose that A1 and A3 hold. Given $x\in\mathbb{R}^{n}$, a set $\left\{z_{i}\right\}_{i=1}^{m}\subset\mathbb{R}^{n}$ of Rademacher random vectors, $s\in\left\{1,\ldots,n\right\}$ and $h>0$, let $y\in\mathbb{R}^{m}$ and $Z\in\mathbb{R}^{m\times n}$ be defined by (3) and (4), respectively. Denote by $\left\{v^{\ell}\right\}_{\ell\geq 0}$ the sequence generated by applying CoSaMP to the corresponding problem (6), with starting point $v^{0}=0$. If $Z$ satisfies the $4s$-RIP for a constant $\delta_{4s}(Z)<0.22665$, then for every $\ell\geq 0$,
$\|v^{\ell}-\nabla f(x)\|_{2}\leq\left[\left(\frac{13.2}{\left(\frac{2}{p}-1\right)^{\frac{1}{2}}}+\frac{11}{\left(\frac{1}{p}-1\right)}\right)s^{-\frac{2-p}{2p}}+\left(\frac{1}{2}\right)^{\ell}\right]\|\nabla f(x)\|_{2}+\dfrac{11Ln}{2}h$ (7)
Proof.
By Lemma 2.4,
$y=Z\left[\nabla f(x)\right]_{(s)}+e,$
where
$e=Z\left(\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right)+hw,$
with $\|w\|_{2}\leq Ln/2$. Thus, by Theorem 5 and Remark 3 in [9] we have
$\|v^{\ell}-\left[\nabla f(x)\right]_{(s)}\|_{2}$$\leq$ $\left(\frac{1}{2}\right)^{\ell}\|\left[\nabla f(x)\right]_{(s)}\|_{2}+11\|e\|_{2}$
$\leq$ $\left(\frac{1}{2}\right)^{\ell}\|\nabla f(x)\|_{2}+11\left\|Z\left(\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right)\right\|_{2}+11h\|w\|$
$\leq$ $\left(\frac{1}{2}\right)^{\ell}\|\nabla f(x)\|_{2}+11\left\|Z\left(\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right)\right\|_{2}+\frac{11Ln}{2}h.$
Then, using Lemma 2.1, we obtain
$\|v^{\ell}-\nabla f(x)\|_{2}$$\leq$ $\|v^{\ell}-\left[\nabla f(x)\right]_{(s)}\|_{2}+\|\left[\nabla f(x)\right]_{(s)}-\nabla f(x)\|_{2}$ (8)
$\leq$ $11\left\|Z\left(\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right)\right\|_{2}+\frac{11Ln}{2}h$
$+\left[\left(\frac{1}{2}\right)^{\ell}+\frac{s^{\frac{1}{2}-\frac{1}{p}}}{\left(\frac{2}{p}-1\right)^{\frac{1}{2}}}\right]\|\nabla f(x)\|_{2}$
From Lemmas 2.1 and 2.2, and the assumption $\delta_{4s}(Z)\leq 0.22665$, we also have
$\left\|Z\left(\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right)\right\|_{2}$$\leq$ $\sqrt{1+\delta_{4s}(Z)}\left(\left\|\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right\|+\frac{1}{\sqrt{s}}\left\|\nabla f(x)-\left[\nabla f(x)\right]_{(s)}\right\|_{1}\right)$ (9)
$\leq$ $\sqrt{1.22665}\left(\frac{1}{\left(\frac{2}{p}-1\right)^{\frac{1}{2}}}+\frac{1}{\left(\frac{1}{p}-1\right)}\right)s^{\frac{1}{2}-\frac{1}{p}}\|\nabla f(x)\|_{2}$
Finally, combining (8) and (9), we conclude that
$\|v^{\ell}-\nabla f(x)\|_{2}\leq\left[\left(\frac{11\sqrt{1.22665}+1}{\left(\frac{2}{p}-1\right)^{\frac{1}{2}}}+\frac{11}{\left(\frac{1}{p}-1\right)}\right)s^{-\frac{2-p}{2p}}+\left(\frac{1}{2}\right)^{\ell}\right]\|\nabla f(x)\|_{2}+\frac{11Ln}{2}h,$
which implies that (7) is true.
The following corollary may be deduced from Lemma 2.5 by direct computation.
Corollary 2.6.
Let $\theta,\epsilon\in(0,1)$. Under the assumptions of Lemma 2.5, if
$\ell=\left\lceil\dfrac{\log(\theta/4)}{\log(0.5)}\right\rceil,$ (10)
$s\geq s(\theta,p):=\left[\frac{4}{\theta}\left(\frac{13.2}{\left(\frac{2}{p}-1\right)^{\frac{1}{2}}}+\frac{11}{\left(\frac{1}{p}-1\right)}\right)\right]^{\frac{2p}{2-p}},$ (11)
and
$0<h\leq\dfrac{\theta}{11Ln}\epsilon,$ (12)
then, for $g=v^{\ell}$ we have
$\|g-\nabla f(x)\|\leq\dfrac{\theta}{2}\|\nabla f(x)\|+\dfrac{\theta}{2}\epsilon.$ (13)
We call the quantity $s(\theta,p)$ defined in (11) the effective sparsity level.
2.2 Sufficient Decrease Condition
From Corollary 2.6, we see that if $s$ is sufficiently large and $h$ is sufficiently small, then CoSaMP applied to (6) is able to find a vector $g$ that satisfies
$\|g-\nabla f(x)\|\leq\dfrac{\theta}{2}\|\nabla f(x)\|+\dfrac{\theta}{2}\epsilon$ (14)
whenever the matrix $Z$ satisfies the $4s$-RIP with parameter $\delta_{4s}(Z)<0.22665$. The next lemma justifies our interest in this type of gradient approximation.
Lemma 2.7.
Suppose that A1 holds. Given $\epsilon\in(0,1)$ and $\theta\in[0,1/2)$, let $x\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{n}$ be such that $\|\nabla f(x)\|>\epsilon$ and (14) hold. Moreover, let
$x^{+}=x-\frac{1}{\sigma}g$
for some $\sigma>0$. If
$\sigma\geq\dfrac{(\theta+1)^{2}}{(1-2\theta)}L$ (15)
then
$f(x)-f(x^{+})\geq\dfrac{1}{2\sigma}\epsilon^{2}.$ (16)
Proof.
Since $\|\nabla f(x)\|>\epsilon$, it follows from (14) that
$\|\nabla f(x)-g\|\leq\theta\|\nabla f(x)\|.$ (17)
Using assumption A1, the Cauchy-Schwarz inequality and (17) we get
$f(x^{+})$$\leq$ $f(x)+\nabla f(x)^{T}\left(-\frac{1}{\sigma}g\right)+\dfrac{L}{2}\|\frac{1}{\sigma}g\|^{2}$ (18)
$=$ $f(x)-\frac{1}{\sigma}\nabla f(x)^{T}g+\dfrac{L}{2\sigma^{2}}\|g\|^{2}$
$=$ $f(x)-\frac{1}{\sigma}\nabla f(x)^{T}\left[\nabla f(x)-\nabla f(x)+g\right]+\dfrac{L}{2\sigma^{2}}\|g\|^{2}$
$\leq$ $f(x)-\frac{1}{\sigma}\|\nabla f(x)\|^{2}+\frac{1}{\sigma}\|\nabla f(x)\|\|\nabla f(x)-g\|+\dfrac{L}{2\sigma^{2}}\|g\|^{2}$
$\leq$ $f(x)-\frac{1}{\sigma}\|\nabla f(x)\|^{2}+\dfrac{\theta}{\sigma}\|\nabla f(x)\|^{2}+\dfrac{L}{2\sigma^{2}}\|g\|^{2}.$
On the other hand, by (14) we also have
$\|g\|\leq\|g-\nabla f(x)\|+\|\nabla f(x)\|\leq\theta\|\nabla f(x)\|+\|\nabla f(x)\|=(\theta+1)\|\nabla f(x)\|,$
which gives
$\|g\|^{2}\leq(\theta+1)^{2}\|\nabla f(x)\|^{2}.$ (19)
Combining (18) and (19) it follows that
$f(x^{+})$$\leq$ $f(x)-\left(\frac{1}{\sigma}-\frac{\theta}{\sigma}-\frac{L}{2\sigma^{2}}(\theta+1)^{2}\right)\|\nabla f(x)\|^{2}$
$=$ $f(x)-\left(1-\theta-\frac{L}{2\sigma}(\theta+1)^{2}\right)\frac{1}{\sigma}\|\nabla f(x)\|^{2},$
and so
$f(x)-f(x^{+})\geq\left(1-\theta-\frac{L}{2\sigma}(\theta+1)^{2}\right)\frac{1}{\sigma}\|\nabla f(x)\|^{2}.$ (20)
By (15) we have
$1-\theta-\frac{L}{2\sigma}(\theta+1)^{2}\geq\dfrac{1}{2}.$ (21)
Finally, combining (20), (21) and using the inequality $\|\nabla f(x)\|>\epsilon$, we obtain (16). ∎
The lemma below suggests a way to construct a gradient approximation $g$ that satisfies (14) with probability one.
Lemma 2.8.
Suppose that A1 holds. Given $x\in\mathbb{R}^{n}$ and $\epsilon>0$, let $g\in\mathbb{R}^{n}$ be given by
$g_{i}=\dfrac{f(x+he_{i})-f(x)}{h},\quad i=1,\ldots,n,$ (22)
with
$0<h\leq\dfrac{2\theta}{L\sqrt{n}}\epsilon.$ (23)
If $\|\nabla f(x)\|>\epsilon$, then $g$ satisfies (14).
Proof.
By (22), A1, (23) and $\epsilon<\|\nabla f(x)\|$, we have
$\|\nabla f(x)-g\|\leq\dfrac{L\sqrt{n}}{2}h\leq\theta\epsilon<\frac{\theta}{2}\|\nabla f(x)\|+\frac{\theta}{2}\epsilon.$
3 New Method and its Complexity
The proposed method is presented below as Algorithm 1. We call it Fully Adaptive ZORO, or ZORO-FA, as the target sparsity ($s$), sampling radius ($h$), step-size ($1/\sigma$), and number of queries ($m$) are all adaptively selected. ZORO-FA includes a safeguard mechanism, ensuring that it will make progress even when $\nabla f(x)$ does not satisfy the compressibility assumption A3.
{mdframed}
Algorithm 1. Fully Adaptive ZORO, or ZORO-FA.
Step 0. Given $x_{0}\in\mathbb{R}^{n}$, $\epsilon\in(0,1)$, $\theta\in(0,1/2)$, $b\geq 1$, and $\sigma_{0}>0$. Choose $s_{0}$ such that $\lceil bs_{0}\ln(n)\rceil\leq n$, sample i.i.d. Rademacher random vectors $\{z_{i}\}_{i=1}^{n}$, and set $k:=0$.
Step 1. Set $j:=0$.
Step 2.1 Set $s_{j}=2^{j}s_{0}$, $\sigma_{j}=2^{j}\sigma_{0}$, and $m_{j}=\left\lceil bs_{j}\ln(n)\right\rceil$. If $m_{j}\geq n$ go to Step 2.4.
Step 2.2 Set
$h_{j}=\dfrac{\theta}{11n\sigma_{j}}\epsilon\quad\text{and}\quad Z^{(j)}=\frac{1}{\sqrt{m_{j}}}\begin{bmatrix}z_{1}&\ldots&z_{m_{j}}\end{bmatrix}.$
Compute $y_{k}^{(j)}\in\mathbb{R}^{m_{j}}$ by
$\left[y^{(j)}_{k}\right]_{i}=\dfrac{f(x_{k}+h_{j}z_{i}^{(j)})-f(x_{k})}{h_{j}\sqrt{m_{j}}},\quad i=1,\ldots,m_{j}.$
Step 2.3 Compute
$g_{k}^{(j)}\approx\arg\min_{\|g\|_{0}\leq s_{j}}\|Z^{(j)}g-y^{(j)}_{k}\|_{2}$
by applying $\left\lceil\dfrac{\log\left(\theta/4\right)}{\log(0.5)}\right\rceil$ iterations of CoSaMP. Go to Step 2.5.
Step 2.4 For
$h_{j}=\frac{2\theta}{\sigma_{j}\sqrt{n}}\epsilon,$
compute $g_{k}^{(j)}\in\mathbb{R}^{n}$ by
$[g_{k}^{(j)}]_{\ell}=\frac{f(x_{k}+h_{j}e_{\ell})-f(x_{k})}{h_{j}},\ \ell=1,\ldots,n.$
Step 2.5 Let $x_{k,j}^{+}=x_{k}-\left(\frac{1}{\sigma_{j}}\right)g_{k}^{(j)}$. If
$f(x_{k})-f(x_{k,j}^{+})\geq\dfrac{1}{2\sigma_{j}}\epsilon^{2}$ (24)
holds, set $j_{k}=j$, $g_{k}=g_{k}^{(j)}$, $s_{k}=s_{j_{k}}$, $\sigma_{k}=\sigma_{j_{k}}$ and go to Step 3. Otherwise, set $j:=j+1$ and go to Step 2.1.
Step 3. Define $x_{k+1}=x_{k,j_{k}}^{+}$ set $k:=k+1$ and go to Step 1.
Remark 3.1.
Note that ZORO-FA bears some resemblence to the adaptive ZORO, or adaZORO algorithm presented in [6]. However, that algorithm only selects the target sparsity ($s$) adaptively. Moreover, we are able to provide theoretical complexity guarantees for ZORO-FA, whereas no such guarantees are presented for adaZORO.
3.1 Oracle Complexity Without Compressible Gradients
Lemma 3.2.
Suppose that A1 holds. If $\|\nabla f(x_{k})\|>\epsilon$, then
$0\leq j_{k}<\max\left\{1,\log_{2}\left(\dfrac{2n}{bs_{0}\ln(n)}\right),\log_{2}\left(\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}.$ (25)
Proof.
From Steps 1 and 2 of Algorithm 1, we see that $j_{k}$ is the smallest non-negative integer for which (24) holds. Suppose by contradiction that (22) is not true, i.e.,
$j_{k}\geq\max\left\{1,\log_{2}\left(\dfrac{2n}{bs_{0}\ln(n)}\right),\log_{2}\left(\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}.$ (26)
Then
$m_{j_{k}-1}=\lceil{bs_{j_{k}-1}\ln(n)\rceil}\geq b\left(2^{j_{k}-1}s_{0}\right)\ln(n)\geq b\left(\dfrac{n}{b\ln(n)}\right)\ln(n)=n.$
Consequently, by Steps 2.1 and 2.4 of Algorithm 1, it follows that $g_{k}^{(j_{k}-1)}$ was defined by
$\left[g_{k}^{(j_{k}-1)}\right]_{i}=\dfrac{f(x_{k}+h_{j_{k}-1}e_{i})-f(x_{k})}{h_{j_{k}-1}},\quad i=1,\ldots,n,$ (27)
with
$h_{j_{k}-1}=\dfrac{2\theta}{\sigma_{j_{k}-1}\sqrt{n}}\epsilon.$ (28)
In addition, (26) also implies that
$\sigma_{j_{k}-1}=2^{j_{k}-1}\sigma_{0}\geq\dfrac{(\theta+1)^{2}}{(1-2\theta)}L$ (29)
Since $\theta\in(0,1)$, (28) and (29) imply that
$0<h_{j_{k}-1}<\dfrac{2\theta}{L\sqrt{n}}\epsilon.$ (30)
In view of (27), (30) and assumption $\|\nabla f(x_{k})\|>\epsilon$, it follows from Lemma 2.8 that (14) holds for $x=x_{k}$ and $g=g_{k}^{(j_{k}-1)}$. Thus, by (29) and Lemma 2.7 we would have
$f(x_{k})-f(x_{k,j_{k}-1}^{+})\geq\dfrac{1}{2\sigma_{j_{k}-1}}\epsilon^{2},$
contradicting the definition of $j_{k}$. ∎
As a consequence of Lemma 3.2, we have the following result.
Lemma 3.3.
Suppose that A1 holds. If $\|\nabla f(x_{k})\|>\epsilon$ then
$s_{k}\leq\max\left\{2s_{0},\left[\dfrac{2n}{b\ln(n)}\right],\left[\dfrac{2(\theta+1)^{2}L}{(1-2\theta)\sigma_{0}}\right]s_{0}\right\}.$
and
$\sigma_{k}\leq\max\left\{2\sigma_{0},\left[\dfrac{2n}{bs_{0}\ln(n)}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}.$
Denote
$T(\epsilon)=\inf\left\{k\in\mathbb{N}\,:\,\|\nabla f(x_{k})\|\leq\epsilon\right\}.$ (31)
Theorem 3.4.
Suppose that A1-A2 hold and let $\left\{x_{k}\right\}_{k=0}^{T(\epsilon)}$ be generated by Algorithm 1. Then
$T(\epsilon)\leq 2\max\left\{2\sigma_{0},\left[\dfrac{2n}{bs_{0}\ln(n)}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}(f(x_{0})-f_{low})\epsilon^{-2}.$ (32)
Proof.
From Step 2.5 of Algorithm 1, we have
$f(x_{k})-f(x_{k+1})\geq\dfrac{1}{2\sigma_{k}}\epsilon^{2},\quad\text{for}\,\,k=0,\ldots,T(\epsilon)-1.$
Summing up these inequalities and using A2 and Lemma 3.3, we obtain
$f(x_{0})-f_{low}$$\geq$ $\sum_{k=0}^{T(\epsilon)-1}f(x_{k})-f(x_{k+1})\geq\sum_{k=0}^{T(\epsilon)-1}\dfrac{1}{2\sigma_{k}}\epsilon^{2}$
$\geq$ $\dfrac{T(\epsilon)}{2\max\left\{2\sigma_{0},\left[\dfrac{2n}{bs_{0}\ln(n)}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}}\epsilon^{2},$
which implies that (32) is true. ∎
Corollary 3.5.
Suppose that A1-A2 hold and denote by $FE_{T(\epsilon)}$ the total number of function evaluations performed by Algorithm 1 up to the $(T(\epsilon)-1)$-st iteration. Then
$FE_{T(\epsilon)}$$\leq$ $2(n+1)\max\left\{1,\log_{2}\left(\dfrac{2n}{bs_{0}\ln(n)}\right),\log_{2}\left(\dfrac{2(\theta+1)^{2}}{1-2\theta}\dfrac{L}{\sigma_{0}}\right)\right\}\times$ (33)
$\left[2\max\left\{2\sigma_{0},\left[\dfrac{2n}{bs_{0}\ln(n)}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}(f(x_{0})-f_{low})\epsilon^{-2}\right].$
Proof.
Let us consider the $k$th iteration of Algorithm 1. For any $j\in\left\{0,\ldots,j_{k}\right\}$, the computation of $g_{k}^{(j)}$ requires at most $n+1$ function evaluations. Moreover, checking the acceptance condition (24) at Step 2.5 requires one additional function evaluation. Thus, the total number of function evaluations at the $k$th iteration is bounded from above by $(n+1)(j_{k}+1)$. Therefore, by Lemma 3.2,
$FE_{T(\epsilon)}$$\leq$ $(n+1)\sum_{k=0}^{T(\epsilon)-1}(j_{k}+1)\leq 2(n+1)\sum_{k=0}^{T(\epsilon)-1}j_{k}$ (34)
$\leq$ $2(n+1)\max\left\{1,\log_{2}\left(\dfrac{2n}{bs_{0}\ln(n)}\right),\log_{2}\left(\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}T(\epsilon).$
Finally, combining (34) and Theorem 3.4 we conclude that (33) is true. ∎
Remark 3.6.
Recall that $m_{0}=bs_{0}\log(n)$. Thus, it follows from Corollary 3.5 that
$FE_{T(\epsilon)}\leq\mathcal{O}\left(n\left[\left(\dfrac{2n}{m_{0}}\right)\log_{2}\left(\dfrac{2n}{m_{0}}\right)\right]\epsilon^{-2}\right).$
Therefore, when $m_{0}=n$, then Algorithm 1 requires no more than $\mathcal{O}\left(n\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of $f(\,\cdot\,)$. Regarding the dependence on $n$ and $\epsilon$, this bound is consistent,in order, with the complexity bounds established in [10, 11]. However, to take advantage of potential compressibility of gradients, one should take $m_{0}<<n$. In this case, the worst-case complexity bound becomes $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$.
3.2 Oracle Complexity Assuming Compressible Gradients
Notice that $J=\left\lceil\log_{2}\left(\frac{n}{bs_{0}\ln(n)}\right)\right\rceil$ is the smallest non-negative integer such that $m_{J}=\lceil bs_{J}\ln(n)\rceil\geq n$. Therefore,
$Z^{(j)}=\dfrac{1}{\sqrt{m_{j}}}\left[z_{1}\ldots z_{m_{j}}\right],\quad j=0,\ldots,J-1$
are the only possible sensing matrices used in Algorithm 1. In this subsection we will consider the following additional assumptions:
  • A4.
    Under A1 and A3, the index
    $j^{*}:=\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil,$ (35)
    with $s(\theta,p)$ defined in (11), belongs to the interval $[1,J-1)$.
  • A5.
    For $j^{*}$ defined in (35), the corresponding matrix $Z^{(j^{*})}$ satisfies the $4s_{j^{*}}$-RIP with parameter $\delta_{4s_{j^{*}}}(Z^{(j^{*})})<0.22665$.
Lemma 3.7.
Suppose that A1, A4 and A5 hold. If $\|\nabla f(x_{k})\|>\epsilon$, then
$0\leq j_{k}\leq\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil.$ (36)
Proof.
Let us show that, for $j^{*}$ defined in (35), the corresponding point $x_{k,j^{*}}^{+}$ satisfies (24). Indeed, by A4 and the definition of $J$, it follows that $g_{k}^{(j^{*})}$ was obtained at Step 2.3 of Algorithm 1, using the sensing matrix $Z^{(j^{*})}$, and the vector $y_{k}^{(j^{*})}$ obtained with
$h_{j^{*}}=\dfrac{\theta}{11n\sigma_{j^{*}}}.$ (37)
In view of (35) and Assumption A5, the matrix $Z^{(j^{*})}$ satisfies the $4s_{j^{*}}$-RIP with
$s_{j^{*}}=2^{j^{*}}s_{0}\geq s(\theta,p).$ (38)
Moreover, (35) also implies that
$\sigma_{j^{*}}=2^{j^{*}}\sigma_{0}\geq\dfrac{(\theta+1)^{2}}{(1-2\theta)}L.$ (39)
Combining (37) and (39), we have that
$h_{j^{*}}<\dfrac{\theta}{11nL}\epsilon.$ (40)
Then, it follows from (38), (40) and Corollary 2.6 that (14) holds for $x=x_{k}$ and $g=g_{k}^{(j^{*})}$. Thus, by (39) and Lemma 2.7 we would have
$f(x_{k})-f(x_{k,j^{*}}^{+})\geq\dfrac{1}{2\sigma_{j^{*}}}\epsilon^{2}.$
Since $j_{k}$ is the smallest non-negative integer $j$ for which (24) holds, we conclude that $j_{k}\leq j^{*}$, and so (36) is true. ∎
Lemma 3.7 implies the following upper bounds on $s_{k}$ and $\sigma_{k}$.
Lemma 3.8.
Suppose that A1, A4 and A5 hold. If $\|\nabla f(x_{k})\|>\epsilon$ then
$s_{k}\leq\max\left\{2s_{0},2s(\theta,p),\left[\dfrac{2(\theta+1)^{2}L}{(1-2\theta)\sigma_{0}}\right]s_{0}\right\}.$
and
$\sigma_{k}\leq\max\left\{2\sigma_{0},\left[\dfrac{2s(\theta,p)}{s_{0}}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}.$
Theorem 3.9.
Suppose that A1-A5 hold and let $\left\{x_{k}\right\}_{k=0}^{T(\epsilon)}$ be generated by Algorithm 1, where $T(\epsilon)$ is defined by (31). Then
$T(\epsilon)\leq 2\max\left\{2\sigma_{0},\left[\dfrac{2s(\theta,p)}{s_{0}}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}(f(x_{0})-f_{low})\epsilon^{-2}.$ (41)
Proof.
As in the proof of Theorem 3.4, we have
$f(x_{0})-f_{low}\geq\dfrac{T(\epsilon)}{2\max_{k=0,\ldots,T(\epsilon)-1}\left\{\sigma_{k}\right\}}$ (42)
Thus, combining (42) and the second inequality in Lemma 3.8 we see that (41) is true. ∎
Corollary 3.10.
Suppose that A1-A5 hold and denote by $FE_{T(\epsilon)}$ the total number of function evaluations performed by Algorithm 1 up to the $(T(\epsilon)-1)$-st iteration. Then
$FE_{T(\epsilon)}$$\leq$ $4\left[1+\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right]\times$ (43)
$\max\left\{2s_{0},2s(\theta,p),\left[\dfrac{2(\theta+1)^{2}L}{(1-2\theta)\sigma_{0}}\right]s_{0}\right\}\ln(n)\times$
$\max\left\{2\sigma_{0},\left[\dfrac{2s(\theta,p)}{s_{0}}\right]\sigma_{0},\left[\dfrac{2(\theta+1)^{2}}{(1-2\theta)}\right]L\right\}(f(x_{0})-f_{low})\epsilon^{-2}.$
Proof.
Let us consider the $k$th iteration of Algorithm 1. By Lemma 3.7 and Assumption A4, we have $m_{j}=bs_{j}\ln(n)<n$ for $j=0,\ldots,j_{k}$. Then, for $j\in\left\{0,\ldots,j_{k}\right\}$, the computation of $g_{k}^{(j)}$ requires at most $(m_{j}+1)$ function evaluations. Moreover, checking the acceptance condition (24) requires one additional function evaluation. Thus, the total number of function evaluations at the $k$th iteration is bounded from above by
$\sum_{j=0}^{j_{k}}(m_{j}+2)$$\leq$ $2\sum_{j=0}^{j_{k}}bs_{j}\ln(n)\leq 2(1+j_{k})s_{j_{k}}\ln(n)$
$\leq$ $2\left[1+\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil\right]\times$
$\max\left\{2s_{0},2s(\theta,p),\left[\dfrac{2(\theta+1)^{2}L}{(1-2\theta)\sigma_{0}}\right]s_{0}\right\}\ln(n).$
Consequently
$FE_{T(\epsilon)}$$\leq$ $2\left[1+\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil\right]\times$ (44)
$\max\left\{2s_{0},2s(\theta,p),\left[\dfrac{2(\theta+1)^{2}L}{(1-2\theta)\sigma_{0}}\right]s_{0}\right\}\ln(n)T(\epsilon).$
Finally, combining (44) and Theorem 3.9 we conclude that (43) is true. ∎
When $s_{0}\leq s(\theta,p)$, it follows from Corollary 3.10 that
$FE_{T(\epsilon)}\leq\mathcal{O}\left(\left[\dfrac{s(\theta,p)}{s_{0}}\right]\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right)s(\theta,p)\ln(n)\epsilon^{-2}\right).$
Assuming further that $s_{0}=\mathcal{O}(s(\theta,p))$, i.e., the initial sparsity $s_{0}$ is not too far from $s(\theta,p)$, this bound reduces to
$FE_{T(\epsilon)}\leq\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right).$
This upper complexity bound relies on the assumption that $Z^{(j^{*})}$ satisfy the $4s_{j^{*}}$-RIP with $\delta_{4s_{j^{*}}}(Z^{(j^{*})})<0.22665$ (Assumption A5). In view of Lemma 2.3 and our construction of matrices $Z^{(j)}$, if A1-A4 hold and constant $b$ is sufficiently large, A5 holds with with high probability. Consequently, with high probability, Algorithm 1 needs no more than $\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of $f(\,\cdot\,)$.
4 Illustrative Numerical Experiments
In the following experiments, we compare ZORO-FA to a representative sample of deterministic DFO algorithms, namely Nelder-Mead [20], DFQRM [11], and ZORO [6]. We avoid methods which incorporate curvature, or “second-order” information such as CARS [12], NEWUOA [21, 27] or the derivative-free variation of L-BFGS studied in [24], noting that such functionality could be added to ZORO-FA fairly easily.
4.1 Sparse Gradient Functions
We test ZORO-FA on a challenging test function known to exhibit gradient sparsity, namely the max-s-squared function ($f_{ms}$) [5, 23] defined as
$f_{ms}(x)=\sum_{i=1}^{s}x_{\pi(i)}^{2}\quad\text{ where }|x_{\pi(1)}|\geq|x_{\pi(2)}|\geq\ldots\geq|x_{\pi(n)}|,$
$\pi(\cdot)$ is the permutation ordering the coordinates of $x$ from largest to smallest in magnitude. This function is challenging as, while $\nabla f_{ms}$ is $p$-compressible (in fact, sparse), the location of the largest-in-magnitude entries $\nabla f_{ms}$ changes frequently. Moreover, $\nabla f_{ms}$ is not Lipschitz continuous. We take $n$, the dimension of the function, to be $1000$ and $s=30$.
Parameters.
We tuned the parameters of each algorithm to the best of our ability; see Table 1 for exact values used. In all trials $x_{0}$ is a random draw from the multivariate normal distribution with mean 00 and covariance $10I$. Each algorithm is given a query budget of $500(n+1)$.
Algorithm $b$ step size $s_{0}\dagger$ $\epsilon$ $\theta\ddagger$ $\sigma_{0}$ $\sigma_{\min}$ min. simplex size
DFQRM - - - $10^{-5}$ 00 $1$ $0.01$
ZORO $1$ $0.5$ $30$
adaZORO $1$ $0.5$ $30$
ZORO-FA $1$ $20$ $0.01$ $0.25$ $1$
Nelder-Mead - - - - - - - $0.001$
Table 1: Parameters used in sparse benchmarking experiments of Section 4.1. $\dagger$: Here $s_{0}$ refers to the target sparsity in ZORO, and the initial target sparsity of adaZORO and ZORO-FA (which both dynamically adjust this parameter). $\ddagger$: In DFQRM, $\theta$ has a slightly different meaning.
Results.
Figure 1 displays a typical trajectory, as well as box plots for the run time over ten trials. Note that the only source of randomness in each trial is the initial point $x_{0}$. Unsurprisingly, ZORO and adaZORO perform well, as they are given the exact sparsity as an input and have hand-tuned step-sizes. While Nelder-Mead and DFQRM struggle, ZORO-FA makes steady progress.
Figure 1: Left: Typical trajectories for all five algorithms. Note the $x$-axis displays the cumulative number of queries divided by $n+1$. Right: Wall-clock run time.
4.2 Moré-Garbow-Hillstrom Test Functions
Next, we benchmark ZORO-FA on a subset of the well-known Moré-Garbow-Hillstrom (MGH) suite of test problems [16].
Functions.
We select the 18 problems with variable dimensions $n$ and considered three values of $n$: $n=100$, $n=500$ and $n=1000$. We synthetically increase the number of functions to 36 by toggling the initial point $x_{0}=10^{s}\tilde{x}_{0}$, for $s\in\{0,1\}$ and $\tilde{x}_{0}$ the default initial point associated with the MGH function in question.
Parameters.
We tuned the parameters of each algorithm to the best of our ability but use the same parameter values for all test functions, for all values of $n$, see Table 2 for precise values.
Algorithm $b$ step size $s_{0}\dagger$ $\epsilon$ $\theta$ $\sigma_{0}$ $\sigma_{\min}$ min. simplex size
DFQRM - - - $10^{-5}$ 00$\ddagger$0 ‡ $1$ $0.01$
ZORO $1$ $0.005$ $0.1n$
adaZORO $1$ $0.005$ $0.1n$
ZORO-FA $1$ $0.1n$ $0.01$ $0.25$ $1$
Nelder-Mead - - - - - - - $0.001$
Table 2: Parameters used in the Moré-Garbow-Hillstrom benchmarking experiments of Section 4.2. $\dagger$: Here $s_{0}$ refers to the target sparsity in ZORO, and the initial target sparsity of adaZORO and ZORO-FA (which both dynamically adjust this parameter). $n$ refers to the problem size. $\ddagger$: In DFQRM, $\theta$ has a slightly different meaning.
Data Profiles.
We present the results as data profiles [17] for various tolerances $\tau$. Let $\mathcal{P}$ denote the test problem set. A point $(\alpha,y)$ on a curve in a data profile indicates that a fraction of $y|\mathcal{P}|$ problems were solved, to tolerance $\tau$, given $(n+1)\times\alpha$ function evaluations. We say an algorithm solves a problem to tolerance $\tau$ if it finds an iterate $x_{k}$ satisfying
$f(x_{k})\leq f_{L}+\tau\left(f(x_{0})-f_{L}\right),$ (45)
where $f_{L}$ is the smallest objective function values attained by any algorithm on the problem.
Results.
The results are shown in Figure 2. We observe that ZORO-FA is quickly able to solve a subset of MGH problems, as shown by the left-hand side of each data profile. We found that adaZORO and ZORO struggled in all settings. We suspect this is because (i) Only some MGH test functions exhibit gradient sparsity, and (ii) The Lipschitz constants, and consequently the ideal step sizes, vary greatly among MGH functions. Thus, algorithms which are able to adaptively select their step size have an innate advantage.
Figure 2: Data profiles for various tolerances ($\tau$) and problem dimensions ($n$). Top Row: $n=100$. Middle Row: $n=500$. Bottom Row: $n=1000$. Left Column: $\tau=10^{-1}$. Middle Column: $\tau=10^{-2}$. Right Column: $\tau=10^{-3}$. In the top row, plots for ZORO and adaZORO are absent as they did not solve any problems to the required tolerance.
4.3 Discussion
ZORO-FA outperforms DFQRM and Nelder-Mead on functions exhibiting gradient sparsity, and is in turn outperformed by ZORO and adaZORO. The latter statement is not surprising, as ZORO and adaZORO are given the exact gradient sparsity parameter and a hand-tuned step-size. The situation is reversed for the MGH test functions; ZORO and adaZORO struggle — presumably as they are unable to adjust their step sizes — while Nelder-Mead, DFQRM, and ZORO-FA perform well. It was surprising, at least to us, that Nelder-Mead performed so well. We note however that this performance came at high computational expense, with Nelder-Mead typically taking 10–100 times longer to run (see the right panel in Figure 1).
4.4 Hidden Gradient Compressibility
In contrast to Section 4.1, the MGH functions of Section 4.2 are not explicitly designed to possess sparse or compressible gradients. Nonetheless we observe that ZORO-FA works well, and is frequently able to make progress with $s_{k}\ll n$. To investigate this further, we plot the magnitudes of the entries of $\nabla f(x)$, sorted from largest to smallest, for several MGH test functions at $20$ randomly selected $x$. Letting $|\nabla f(x)|_{(j)}$ denote the $j$-th largest-in-magnitude component of $\nabla f(x)$, we plot the mean of $|\nabla f(x)|_{(j)}$ over all $20$ values of $x$, as well as the minimum and maximum values of $|\nabla f(x)|_{(j)}$ observed. The results are displayed in Figure 3. As is clear, some MGH test functions (e.g., rosex) exhibit gradient compressibility, while others (e.g., trig) do not. This further emphasizes the practical need for optimization methods which can take advantage of the presence of gradient sparsity, yet are robust to its absence.
Figure 3: Gradient magnitude profiles for two MGH functions: rosex (Left), and trig (Right). The solid line represents the mean (over 20 trials), while the shading represents the minimum and maximum values of gradient component magnitudes observed.
5 Conclusion
In this paper, we presented a fully adaptive variant of ZORO, originally proposed in [6], for derivative-free minimization of a smooth functions with $n$ variables and $L$-Lipschitz continuous gradient. The new method, called ZORO-FA, does not require the knowledge of the Lispchitz constant $L$ or the effective sparsity level (when the gradients are compressible). At each iteration, ZORO-FA attempts to exploit a possible compressibility of gradient to compute a suitable gradient approximation using fewer than $\mathcal{O}\left(n\right)$ function evaluations. When the corresponding trial points fail to ensure a functional decrease of $\mathcal{O}\left(\epsilon^{2}\right)$, the method computes gradient approximations via forward finite-differences, which guarantee that the trial points will yield the desired functional decrease. With this safeguard procedure, we proved that, with probability one, ZORO-FA needs no more than $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of the objective function. When the gradient vectors of the objective function are $p$-compressible for some $p\in(0,1)$, we proved that, with high probability, ZORO-FA has an improved worst-case complexity of $\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right)$ function evaluations, where $\theta\in(0,1/2)$ is a user-defined parameter that controls the relative error of the gradient approximations, and $s(\theta,p)$ is the effective sparsity level as defined in Corollary 2.6. Our preliminary numerical results indicate that ZORO-FA is able to exploit the presence of gradient compressibility, while being robust towards its absence.